40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 22 02:43:28 2011, kiwamu, Update, General, in-vac work 
    Reply  Mon Aug 22 21:32:10 2011, kiwamu, Update, General, more in-vac work : AS clipping fixed and OSEM/oplev adjustment 
       Reply  Tue Aug 23 06:49:24 2011, Anamaria, Update, General, more in-vac work : AS clipping fixed and OSEM/oplev adjustment 
          Reply  Tue Aug 23 09:40:37 2011, kiwamu, Update, General, Re: AS clipping fixed 
Message ID: 5279     Entry time: Mon Aug 22 21:32:10 2011     In reply to: 5275     Reply to this: 5284
Author: kiwamu 
Type: Update 
Category: General 
Subject: more in-vac work : AS clipping fixed and OSEM/oplev adjustment 

[Keiko / Jenne / Jamie / Kiwamu]

 We did the following things today :

  + fixed the AS clipping issue

  + realigned all the oplevs

  + checked and adjusted the all OSEM DC values, including PRM, SRM, BS, ITMs, ETMs, MC1 and MC3

 

Since we touched the OSEMs the alignment has changed somewhat.

Right now Jenne, Suresh and I are working on the "confirmation alignment".

Once we find the alignment is still good (steerable by the PZTs and the DC coil bias), tomorrow we will do the drag&wipe and door closing.

Quote from #5275

We need to check/fix the AS beam clipping and once it's done we will readjust the OSEM mid range and the oplevs.

 

ELOG V3.1.3-